1. How many non-bonding electrons does phosphorus have?




2. Complete and balance this reaction: sulfuric acid reacts with barium chloride



3. Identify the precipitate in this reaction: calcium nitrate reacts with sodium phosphate.
1
Expert's answer
2016-06-16T08:32:02-0400
1. How many non-bonding electrons does phosphorus have?
Phosphorus has [Ne]3s23p3 electron configuration. That’s why it has five (2+3=5) non-bonding electrons.
2. Complete and balance this reaction: sulfuric acid reacts with barium chloride
H2SO4 + BaCl2 = BaSO4↓ + 2HCl
3. Identify the precipitate in this reaction: calcium nitrate reacts with sodium phosphate.
3Ca(NO3)2 + 2Na3PO4 = Ca3(PO4)2↓ + 6NaNO3